Experiment - Enthalpy of Neutralization

Use Hess' Law to Calculate the enthalpy of dissociation for chloroacetic acid.

H (aq) + OH -> H2O (l) ΔHn(HCl)

HC2H2ClO2(aq) + OH (aq) -> C2H2ClO2 + H2O (l) ΔHn(HC2H2ClO2)

We used a strong acid, HCl as the limiting reactant and NaOH as the the excess.

Weak acid: HC2H2ClO2 (aq) + H (aq) + C2H2ClO2 (aq) ΔHd(HC2H2ClO2)

HCl: Average ΔHn = -59 KJ

HC2H2ClO2: Average ΔHn = -61 KJ
Answer #1

Eq.1; H+ (aq) + OH-(aq) H2O (l) ; 1 = -59 KJ

Eq.2 ; HC2H2ClO2 (aq) + OH- (aq) C2H2ClO2 + H2O ; 2= -61 KJ

Eq. 3 ; HC2H2ClO2 (aq) H+ (aq) + C2H2ClO2 (aq) .

Now,

Eq.3 = Eq.2 - Eq.1

So, by Hess law

= 2 - 1

= {- 61 - ( -59) } KJ

= - 2 KJ.

So, enthalpy of dissociation of chloroacetic acid = -2 KJ.